Huntington disease is a form of neuromuscular degeneration that results from a dominant mutation on the autosome. The early onset reduces the reproductive fitness of a affected individuals by about 20%. A large study in Michigan estimated that the frequency of the dominant allele is about 5 x10-6. Assume that this population undergoes random mating and is in equilibrium. What is the expected frequency of individuals who carry the dominant allele? What is the estimated mutation rate of the dominant deleterious allele?
Answer #1

The frequency of dominant allele is 5 x 10-6

The population is random mating and is in equilibrium according to which p+ q = 1

p is the frequency of dominant allele and q is the frequency or recessive allele.

p2 + q2 + 2pq = 1

The expected frequency of people who carry the dominant allele is p2 + 2pq , where p2 is the frequency of homozygous dominant individuals and 2pq is the frequency of heterozygous individuals.

p2 = (5 x 10-6 )2 = 25 x 10-12

p + q = 1

5 x 10-6 + q = 1

q = 1- 5 x 10-6 = 0.999995

2pq = 2 x 5 x 10-6 x 0.999995 = 9.9 x 10-6

25 x 10-12 + 9.9 x 10-6 = 0.000025 x 10-6 + 9.9 x 10-6 = 9.9 x 10-6

The expected frequency of individuals carrying the dominant allele is 9.9 x 10-6